OpenStudy (anonymous):

I Will Fan You***** Given the function f(x) = 5x, Section A is from x = 0 to x = 1 and Section B is from x = 2 to x = 3. Part A: Find the average rate of change of each section. (4 points) Part B: How many times greater is the average rate of change of Section B than Section A? Explain why one rate of change is greater than the other. (6 points)

OpenStudy (anonymous):

@Jadeishere

OpenStudy (jadeishere):

i don't understand the question

OpenStudy (princeharryyy):

A))

OpenStudy (princeharryyy):

A))) f(0) = 5*0 =0 f(1) = 5*1 =5 Average rate = f(1) -f(0) / (1-0) = 5-0 / 1-0 =5

OpenStudy (princeharryyy):

that's for section A, Similarly for section b, f(2) = 5*2 =10 f(3) =5*3 =15 f(3) -f(2) / 3-2 = 15-10 / 3-2 =5

OpenStudy (princeharryyy):

B)) since the function in a linear function, which means a straight line, the average rate is same for both the given sections for section a and section b.